7.2 Jay-Dubs Disc Golf

Jonathan has a passion for disc golf. About 10 years ago, his very good friend Andy introduced him to the sport. Since then, the two of them have become masters of the sport. The two play several times a week and travel across the country seeking out new courses to practice their skills. One evening, while sitting around having dinner with their wives, the two men decided that they wanted to quit their jobs and open their own disc golf course. As the night went on, Jonathan’s wife Sara (who is a chartered professional accountant) and Andy’s wife Camille played along with the conversation, thinking that it was all an unrealistic dream. About a week later, the two men called a meeting together with their wives. They were serious about the idea and had developed a plan to make their dream become a reality. Below is an account of the conversation: Andy: Ladies, we have brought you here today to discuss our plans to start our very own disc golf course,

“Jay-Dubs Disc Golf.” We really think that this is going to work, but will need your support in order to make it happen.

Camille: Okay, let’s hear it. Andy: Jonathan and I have found the perfect place for the course. It is a large green piece of land that we are able to rent in Milton, Ontario. It is the perfect place for a course, because not only does it have a growing population, it is close to many larger cities, such as Mississauga, Toronto, Oakville, and Guelph. We spoke to the owners of the land, and they will make us pay rent only during the warmer months—from April to September. The rent they charge is $5,000 per month.

Sara: Okay, what about all the other costs?

Andy: I thought you might ask me that. We have drawn up a spreadsheet with all of the costs associated with the course (Exhibit 1).

Camille: How are you going to make money? All I see is a list of costs here. Every time we have gone to play disc golf, it has been free!

Jonathan: Good point; we are going to charge a fee for the course. We think that people will pay because there aren’t really any disc golf courses in the area. The fee won’t be much. We were thinking about $20 for 18 holes and $10 for 9 holes.

Sara: I don’t mean to go all “accountant” on you, but it seems like you are going to need to sell a lot of rounds of golf in order to cover these costs. Also, what about making money to live?!

Jonathan: I knew you would ask about that, Sara. Andy and I will both quit our jobs and will start a painting business. When I was in college, I painted during the summers. It was great money and I loved it. We will paint during the warmer months part-time, and when it is cold and the course is closed, we will paint full-time. Don’t worry, I know what you are going to ask next. We have analyzed the revenues and expenses associated with painting (Exhibit 2). Based on my previous painting experience, I know that the average painting job is 1,300 square feet. On an annual basis, we are forecasting that for every round of disc golf we sell, we should sell three painting jobs.

Sara: Okay, so how many rounds of disc golf and painting jobs do you need to sell in order to cover your fixed costs?

Andy: Oh, well, I guess we haven’t got there yet. I’m not even sure how to figure that out, to be honest.

Camille: I’m concerned! Sara, to me it isn’t just about covering the fixed costs; I’m worried about them making enough money for us to survive. I don’t think I can support this unless Andy makes at least his current net salary of $60,000 per year.

Sara: Oh, good point, Camille, we need the equivalent of Jonathan’s salary to survive too. He makes a net salary of $60,000 as well.

Required

Assume the role of Sara and prepare a report for Jonathan and Andy showing the level of sales required to cover their fixed costs (break-even) in both businesses and the level of sales required to replace their existing incomes from both businesses. Assume a tax rate of 30%. Your report should include any qualitative factors that Jonathan and Andy should consider in their decision.

EXHIBIT 1 – COSTS ASSOCIATED WITH THE DISC GOLF BUSINESS

Cost

Amount

Comments

Land rent

$5,000 per month

Rent payments will be made in April to September

Equipment rent

$9,000 annually

Property maintenance

$6,000 annually

EXHIBIT 2 – REVENUES AND EXPENSES ASSOCIATED WITH THE PAINTING BUSINESS Revenues/Expenses

Revenues/Expenses

Amount

Comments

Price charged per square foot

$2.00

Paint per square foot

0.32

Rollers (each)

1.49

4 rollers per job

Paint brushes

250.00

Annual cost

Other supplies

800.00

Annual cost

Depreciation - Vehicle

10,000.00

Annual cost

Advertising costs

2,000.00

Annual cost
Answer #1

For the business of Disc Golf;

As the fixed cost = $45,000

And the tax component is 30%, the required revenue before tax would be

(45000 x 30/70) + 45,000 = $64,286(approx.)

And so the revenue to be generated to cover the fixed costs (break even) where there is nil profit after tax will be $64,286.

No. of holes to be sold for break even =64,286÷10/9 or 64,286÷20/18

Which = 58,442 holes (approx. of 58,441.81)

And for an income after which the profit after tax = $60,000

Can be calculated as follows,

Profit after tax [A]             = $60,000

Tax (30%) [B]                      = $25,714 (60,000 x 30/70)

Profit before tax[C=A+B] = $85,714

(+)Fixed cost [D]                = $45,000

So, Revenue [C+D]            = $1,30,714

So the revenue required for a profit after tax = $ 1,30,714

No. of holes to be sold for a revenue of $1,30,714 = 1,30,714÷10/9 (or) 1,30,714÷20/18

=1,18,831 holes (approx. of 1,18,830.90)

For the business of painting;

As the fixed cost = $13,050

And the tax component is 30%, the required revenue before tax would be

13,050 x 30/70 + 13,050 = $18,643(approx.)

And so the revenue to be generated to cover the fixed costs (break even) where there is nil profit after tax will be $18,643.

No. of jobs for break even is as follows,

Revenue for each job                 = $2,600 (1300 x 2)

(-)Expense for paint per sq. foot =$416 (0.32 x 1300)

(-)Expense of rollers for each job = $5.96(1.49 x 4)

Profit from a job of painting       = $2,178.04

So, the number of jobs = 18,643÷2,178.04

Which                              = 9 jobs (approx. of 8.55)

And for an income after which the profit after tax = $60,000

Can be calculated as follows,

Profit after tax [A]             = $60,000

Tax (30%) [B]                      = $25,714 (60,000x30/70)

Profit before tax[C=A+B] =$85,714

(+)Fixed cost [D]                = $13,050

So, Revenue [C+D]            = $ 98,964

So the revenue required for a profit after tax = $ 98,964

No. of holes to be sold for a revenue of $ 98,964 =98,964÷2,178.04

=45 jobs (approx. of 45.43)